# Top-Down Semantic Refinement for Image Captioning

### v1

@@ Introduction
At the intersection of computer vision and natural language processing, Large Vision-Language Models (VLMs)(Radford et al. 2021; Jia et al. 2021; Li et al. 2019; Tan and Bansal 2019; Chen et al. 2023a; Vaswani et al. 2017) have become the dominant force in image captioning. Through powerful visual encoders and language decoders, these models can generate fluent text that is generally aligned with the image content(Li et al. 2022, 2023a; Wang et al. 2022; Radford et al. 2021; Zhang et al. 2025b). However, despite their remarkable success, the core auto-regressive generation mechanism of VLMs exposes a fundamental flaw, i.e., an inherent lack of planning capability. When generating each token, VLMs typically employ greedy or beam search strategies(Brown et al. 2020; Hou et al. 2025; Meister, Vieira, and Cotterell 2021; Radford et al. 2019; Rennie et al. 2017; Zhang et al. 2025e). This decision-making process is inherently “myopic", confined to maximizing local probabilities without “deliberate thought" or foresight and planning capability for the global narrative structure.
@@
This lack of planning capability leads to an intractable dilemma: the model either produces a coherent but detail-poor “safe" description to ensure consistency, or it generates factual errors and logical breaks, i.e., the “hallucination" phenomenon, when attempting to capture rich details without global guidance (Jia et al. 2021; Radford et al. 2021; Rohrbach et al. 2019; OpenAI 2023; Zhang et al. 2025d). To address this challenge, the research community once turned to a seemingly intuitive “bottom-up" paradigm(Anderson et al. 2018; Yang et al. 2017; Zhang et al. 2021; Stefanini et al. 2021; Herdade et al. 2020). These methods first detect independent regions in an image, describe them separately, and finally “stitch" these fragmented descriptions into a complete caption. However, this “local-to-global" strategy fails to address the core problem. Lacking a unified global plan as an anchor from the outset, the resulting descriptions often degenerate into a simple list of facts, leading to semantic fragmentation and logical incoherence(Rohrbach et al. 2019; Alikhani et al. 2020; Bugliarello and Elliott 2021; Zohourianshahzadi and Kalita 2021; Zhang et al. 2025f). This proves that merely stitching details together cannot effectively compensate for the VLM’s lack of planning ability.
@@
We argue that the root of the problem lies in the generation paradigm itself, and the solution is to fundamentally reframe image captioning as a planning problem. To this end, we propose an innovative “top-down" semantic refinement framework (TDSR), which redefines the task from a unidirectional generation process into a coarse-to-fine, goal-oriented, hierarchical planning process(Yao et al. 2019; Yarats and Lewis 2018; Zhang et al. 2025g). This idea, illustrated in Figure 1, mimics the human cognitive process(Marr 1982; Bar and Bubic 2013; Mefford et al. 2023): first, form a holistic impression of the image to generate a high-level, core description as a “planning blueprint" (e.g., for a picture of people playing cards, an initial description might be “a group of people are sitting in a room doing something"). Then, using this blueprint as a guide, purposefully and progressively explore and fill in key details (e.g., further specifying it as “a group of men are sitting around a table, engaged in a game of Texas Hold’em poker," and adding that “on the green felt tabletop lie three community cards and a collection of poker chips"). This “global guidance, local refinement" mechanism ensures that all details serve a unified narrative goal, fundamentally guaranteeing high coherence and richness in the description.
@@
Translating this elegant planning concept into an effective computational process hinges on efficient search and planning within the vast language space(Hoang, Haffari, and Cohn 2017; Wiher, Meister, and Cotterell 2022). We rigorously formalize this process as a Markov Decision Process (MDP(Puterman 1994)) and employ Monte Carlo Tree Search (MCTS(Świechowski et al. 2022; Kemmerling, Lütticke, and Schmitt 2023)) as the core engine. However, directly applying standard MCTS to a VLM is computationally infeasible due to the model’s massive inference cost(Browne et al. 2012a). Therefore, our core technical contribution lies in deeply optimizing the MCTS algorithm to enable efficient planning within VLMs. By incorporating a Visual-Guided Parallel Expansion mechanism and a lightweight value network, our algorithm reduces the call frequency to the expensive VLM by an order of magnitude without sacrificing planning quality, successfully resolving the efficiency bottleneck. Our framework (TDSR), as a plug-and-play module, significantly enhances the performance of existing VLMs and achieves state-of-the-art or highly competitive results on multiple benchmarks.
@@
As a cross-disciplinary field between computer vision and natural language processing(Zhang et al. 2025c; Radford et al. 2021; Jia et al. 2021; Chen et al. 2023a; Zhang et al. 2025a), early research in image captioning primarily adopted the encoder-decoder framework. Methods represented by Show and Tell (Vinyals et al. 2015) used a CNN(Krizhevsky, Sutskever, and Hinton 2012; Lecun et al. 1998) to extract global image features and an RNN(Elman 1990) to generate a fluent description. Subsequently, Show, Attend and Tell (Vinyals et al. 2015) introduced an attention mechanism, allowing the model to focus on local regions of the image. These pioneering works excelled at generating grammatically coherent sentences, but their mechanism of generating a single, holistic description meant they often overlooked fine-grained object details, leading to a lack of richness and specificity. This is precisely one of the core problems our TDSR framework aims to solve through multi-step refinement.
@@
To enhance detail capture, subsequent research shifted towards the “bottom-up" paradigm(Anderson et al. 2018; Yang et al. 2017; Zhang et al. 2021; Stefanini et al. 2021; Herdade et al. 2020). This approach typically first identifies independent objects or regions in an image, generates local descriptions for them, and finally stitches these segments into a complete sentence. DenseCap (Johnson, Karpathy, and Fei-Fei 2015) is a typical representative of this line of work, which utilizes an object detector to locate and describe regions individually. Follow-up works, such as Patch Matters (Peng et al. 2025b) and FineCaption (Hua et al. 2024b), focused on improving the quality of these local descriptions. Although these methods significantly increased detail richness, their “split-first, stitch-later" process inherently decouples from the global context, often leading to semantic fragmentation and insufficient global coherence. In stark contrast, our “top-down" approach, guided by global context, fundamentally avoids the inconsistency issues inherent in the “split-first, stitch-later" process.
@@
In recent years, Large Vision-Language Models (VLMs), such as LLaVA-1.5 (Liu et al. 2023b), Qwen-VL (Bai et al. 2023), and Ferret (You et al. 2023a), have significantly advanced visual narrative capabilities through pre-training on massive image-text data. However, despite their powerful foundational abilities, their standard autoregressive generation still faces the inherent trade-off between detail and coherence. Consequently, several training-free enhancement methods have emerged. For instance, some works employ iterative prompting (e.g., IT (Zhou et al. 2023)) to induce the model to output more details.
@@
To systematically evaluate the generalization and practical effectiveness of TDSR across different model architectures, we deploy it on two widely adopted multimodal large language models: Qwen2.5-VL(Bai et al. 2025) and LLaVA-1.5 (7B)(Liu et al. 2023a). We compare it against a diverse set of representative baselines, which fall into two major paradigms: Training-free image captioning enhancement methods: including IT(Pi et al. 2024), Patch Matters(Peng et al. 2025a), and FINECAPTION, which improve visual description quality without additional training. All baselines in this category are implemented on top of LLaVA-1.5 (7B). Foundation vision-language models: including Shikra-13B(Chen et al. 2023c), MiniGPT-v2(Chen et al. 2023b), Ferret-13B(You et al. 2023b), VisionLLM-H(Wang et al. 2023), KOSMOS-2(Peng et al. 2023), Alpha-CLIP-13B(Sun et al. 2023), and VistaLLM(Pramanick et al. 2023), representing the dominant modeling paradigms in the open-source multimodal field.

### v2

@@ Introduction
At the intersection of computer vision and natural language processing, Large Vision-Language Models (VLMs)(Radford et al. 2021; Jia et al. 2021; Li et al. 2019; Tan and Bansal 2019; Chen et al. 2023a; Vaswani et al. 2017; Chen et al. 2025; Ma et al. 2022; Zhang et al. 2026b) have become the dominant force in image captioning. Through powerful visual encoders and language decoders, these models can generate fluent text that is generally aligned with the image content(Li et al. 2022, 2023a; Wang et al. 2022; Radford et al. 2021; Zhang et al. 2025b). However, despite their remarkable success, the core auto-regressive generation mechanism of VLMs exposes a fundamental flaw, i.e., an inherent lack of planning capability. When generating each token, VLMs typically employ greedy or beam search strategies(Brown et al. 2020; Hou et al. 2025; Meister et al. 2021; Radford et al. 2019; Rennie et al. 2017; Zhang et al. 2025f). This decision-making process is inherently “myopic", confined to maximizing local probabilities without “deliberate thought" or foresight and planning capability for the global narrative structure.
@@
This lack of planning capability leads to an intractable dilemma: the model either produces a coherent but detail-poor “safe" description to ensure consistency, or it generates factual errors and logical breaks, i.e., the “hallucination" phenomenon, when attempting to capture rich details without global guidance (Jia et al. 2021; Radford et al. 2021; Rohrbach et al. 2019; OpenAI 2023; Zhang et al. 2025e; Li et al. 2021). To address this challenge, the research community once turned to a seemingly intuitive “bottom-up" paradigm(Anderson et al. 2018; Yang et al. 2017; Zhang et al. 2021; Stefanini et al. 2021; Herdade et al. 2020). These methods first detect independent regions in an image, describe them separately, and finally “stitch" these fragmented descriptions into a complete caption. However, this “local-to-global" strategy fails to address the core problem. Lacking a unified global plan as an anchor from the outset, the resulting descriptions often degenerate into a simple list of facts, leading to semantic fragmentation and logical incoherence(Rohrbach et al. 2019; Alikhani et al. 2020; Bugliarello and Elliott 2021; Zohourianshahzadi and Kalita 2021; Zhang et al. 2025g). This proves that merely stitching details together cannot effectively compensate for the VLM’s lack of planning ability.
@@
We argue that the root of the problem lies in the generation paradigm itself, and the solution is to fundamentally reframe image captioning as a planning problem. To this end, we propose an innovative “top-down" semantic refinement framework (TDSR), which redefines the task from a unidirectional generation process into a coarse-to-fine, goal-oriented, hierarchical planning process(Yao et al. 2019; Yarats and Lewis 2018; Zhang et al. 2025i). This idea, illustrated in Figure 1, mimics the human cognitive process(Marr 1982; Bar and Bubic 2013; Mefford et al. 2023): first, form a holistic impression of the image to generate a high-level, core description as a “planning blueprint" (e.g., for a picture of people playing cards, an initial description might be “a group of people are sitting in a room doing something"). Then, using this blueprint as a guide, purposefully and progressively explore and fill in key details (e.g., further specifying it as “a group of men are sitting around a table, engaged in a game of Texas Hold’em poker," and adding that “on the green felt tabletop lie three community cards and a collection of poker chips"). This “global guidance, local refinement" mechanism ensures that all details serve a unified narrative goal, fundamentally guaranteeing high coherence and richness in the description.
@@
Translating this elegant planning concept into an effective computational process hinges on efficient search and planning within the vast language space(Hoang et al. 2017; Wiher et al. 2022). We rigorously formalize this process as a Markov Decision Process (MDP(Puterman 1994)) and employ Monte Carlo Tree Search (MCTS(Świechowski et al. 2022; Kemmerling et al. 2023)) as the core engine. However, directly applying standard MCTS to a VLM is computationally infeasible due to the model’s massive inference cost(Browne et al. 2012a). Therefore, our core technical contribution lies in deeply optimizing the MCTS algorithm to enable efficient planning within VLMs. By incorporating a Visual-Guided Parallel Expansion mechanism and a lightweight value network, our algorithm reduces the call frequency to the expensive VLM by an order of magnitude without sacrificing planning quality, successfully resolving the efficiency bottleneck. Our framework (TDSR), as a plug-and-play module, significantly enhances the performance of existing VLMs and achieves state-of-the-art or highly competitive results on multiple benchmarks.
@@
As a cross-disciplinary field between computer vision and natural language processing(Zhang et al. 2025c; Radford et al. 2021; Jia et al. 2021; Chen et al. 2023a; Zhang et al. 2025a, 2026c), early research in image captioning primarily adopted the encoder-decoder framework. Methods represented by Show and Tell (Vinyals et al. 2015; Zhang et al. 2025d, h) used a CNN(Krizhevsky et al. 2012; Lecun et al. 1998; Ma et al. 2024a, b, 2025e, 2025d) to extract global image features and an RNN(Elman 1990; Ma et al. 2025a, b, c; Shen et al. 2025; Long et al. 2025) to generate a fluent description. Subsequently, Show, Attend and Tell (Vinyals et al. 2015) introduced an attention mechanism, allowing the model to focus on local regions of the image. These pioneering works excelled at generating grammatically coherent sentences, but their mechanism of generating a single, holistic description meant they often overlooked fine-grained object details, leading to a lack of richness and specificity. This is precisely one of the core problems our TDSR framework aims to solve through multi-step refinement.
@@
To enhance detail capture, subsequent research shifted towards the “bottom-up" paradigm(Anderson et al. 2018; Yang et al. 2017; Zhang et al. 2021; Stefanini et al. 2021; Herdade et al. 2020). This approach typically first identifies independent objects or regions in an image, generates local descriptions for them, and finally stitches these segments into a complete sentence. DenseCap (Johnson et al. 2015) is a typical representative of this line of work, which utilizes an object detector to locate and describe regions individually. Follow-up works, such as Patch Matters (Peng et al. 2025a; Zhang et al. 2026d, a) and FineCaption (Hua et al. 2024b), focused on improving the quality of these local descriptions. Although these methods significantly increased detail richness, their “split-first, stitch-later" process inherently decouples from the global context, often leading to semantic fragmentation and insufficient global coherence. In stark contrast, our “top-down" approach, guided by global context, fundamentally avoids the inconsistency issues inherent in the “split-first, stitch-later" process.
@@
In recent years, Large Vision-Language Models (VLMs), such as LLaVA-1.5 (Liu et al. 2023b), Qwen-VL (Bai et al. 2023), and Ferret (You and et al. 2023), have significantly advanced visual narrative capabilities through pre-training on massive image-text data. However, despite their powerful foundational abilities, their standard autoregressive generation still faces the inherent trade-off between detail and coherence. Consequently, several training-free enhancement methods have emerged. For instance, some works employ iterative prompting (e.g., IT (Zhou et al. 2023)) to induce the model to output more details.
@@
To systematically evaluate the generalization and practical effectiveness of TDSR across different model architectures, we deploy it on two widely adopted multimodal large language models: Qwen2.5-VL(Bai et al. 2025) and LLaVA-1.5 (7B)(Liu et al. 2023a). We compare it against a diverse set of representative baselines, which fall into two major paradigms: Training-free image captioning enhancement methods: including IT(Pi et al. 2024), Patch Matters(Peng et al. 2025b), and FINECAPTION, which improve visual description quality without additional training. All baselines in this category are implemented on top of LLaVA-1.5 (7B). Foundation vision-language models: including Shikra-13B(Chen et al. 2023c), MiniGPT-v2(Chen et al. 2023b), Ferret-13B(You et al. 2023), VisionLLM-H(Wang et al. 2023), KOSMOS-2(Peng et al. 2023), Alpha-CLIP-13B(Sun et al. 2023), and VistaLLM(Pramanick et al. 2023), representing the dominant modeling paradigms in the open-source multimodal field.